the graph of which equation has an amplitude of 2 and period of pi?
(1) y=2cosx
(2) y=1/2sin2x
(3) y=2cos1/2x
(4) y=-2cos2x

Answer by lwsshak3(11628)

the graph of which equation has an amplitude of 2 and period of pi?
(1) y=2cosx
(2) y=1/2sin2x
(3) y=2cos1/2x
(4) y=-2cos2x
**
Equations for sin and cos functions:
y=Asin(Bx-C), A=amplitude, period=2π/B, phase shift=C/B
y=Acos(Bx-C), A=amplitude, period=2π/B, phase shift=C/B
ans (4) is the correct answer:
amplitude=2 (although function reflected around x-axis)
B=2
period=2π/2=π